OpenStudy (raffle_snaffle):

Nitrogen is used to inflate the tire of an automobile. The volume of the tire is 0.5 m3. The pressure of the surroundings is 100 kPa. What mass, in kg, of nitrogen is required to achieve a gage pressure inside the tire of 180 kPa at 20°C?

OpenStudy (raffle_snaffle):

I have an answer just want to make sure it's right.

OpenStudy (raffle_snaffle):

@zepdrix

OpenStudy (unklerhaukus):

\[pV = nRT=\frac mMRT\\ m = MpV/RT\]

OpenStudy (unklerhaukus):

\[M=14.01\ \text{g/mol}\\ P=80\ \text{kPa}\\ V=0.5\ \text{m}^3\\ T=20\ \text{°C}=\big(273.15+20\big)\text K=293.15\ \text K\\ R=8.3145\ \frac{\text{m}^3\cdot\text{Pa}}{\text K\cdot\text{mol}}\]

OpenStudy (unklerhaukus):

or is \(P=260\ \text{kPa}\)?

OpenStudy (unklerhaukus):

*280\[P = P_\text{gague}+P_\text{ambient}=\big(180+100\big)\ \text{kPa}=280\ \text{kPa}\]

OpenStudy (raffle_snaffle):

we add P gage and P surroundings together?

OpenStudy (raffle_snaffle):

@PRAETORIAN.10

OpenStudy (raffle_snaffle):

my solution is pretty much the same except I subtracted P gage from P surroundings